# Semiparametric KSD test: unifying score and distance-based approaches for goodness-of-fit testing

### v1

@@ 2.2 Nonparametric score-based tests via characteristic IPMs
Tests based on IPMs can be powerful against nonparametric alternatives when sufficiently rich function classes $\mathcal{F}$ are used. The form and expressiveness of $\mathcal{F}$ fundamentally determine the capacity of the aggregated score statistic to capture discrepancies between the null and alternative. This insight has motivated a rich body of literature on hypothesis testing using IPMs [gretton2012kernel, sriperumbudur2012empirical, sejdinovic2013equivalence, liu2016kernelized, paik2023maximum]. In particular, we define an IPM is characteristic if
@@
Examples of popular characteristic IPMs and related divergences include $\{f:f=\mathbbm{1}(A),A\subseteq\mathcal{X}\}$, Kolmogorov-Smirnov (K-S) distance; $\{f:|f(x)-f(y)|/\|x-y\|_{2}\leq 1\}$, Wasserstein-1 (W-1) distance; a reproducing kernel Hilbert space (RKHS): MMD. Through the equivalence in (4), tests based on these IPMs can be interpreted as nonparametric score-based tests under the exponentially tilted models (2), provided the IPM is characteristic. We formally define such tests below:
@@
According to the definition, the simplest way to verify that a score-based test is nonparametric is to confirm that the corresponding IPM is characteristic, as per definition (5). It is known that the K-S and W-1 distances are characteristic under mild assumptions. Similarly, MMD is characteristic when appropriate kernels and regularity conditions are satisfied [gretton2012kernel, muandet2017kernel]. Despite being characteristic, however, critical computation challenges arise in the case of a nonparametric score-based test, when $\mathcal{F}$ is chosen to be sufficiently rich and complex. To illustrate this, we discuss the computational complexity of several test statistics induced by different IPMs below.
@@
In the univariate case ($d=1$), the classical Kolmogorov-Smirnov (K-S) distance between an empirical distribution and a known distribution can be computed efficiently using sorting algorithms, yielding a computational complexity of $O(n\log n)$. However, in the bivariate case ($d=2$), the worst-case computational complexity increases to $O(n^{2})$ [peacock1983two, fasano1987multidimensional]. When $d>2$, computing the K-S distance becomes an NP-hard to compute [gnewuch2009finding].
@@
Similarly, when $d=1$, the W-1 distance can be computed efficiently in $O(n\log n)$ time via sorting. For $d>1$, however, the problem can be formulated as an assignment problem through discretization, and the computational complexity generally grows to $O(n^{3})$ [kuhn1955hungarian], although approximation algorithms exist to improve efficiency [altschuler2017near].
@@
Among nonparametric score-based tests, MMD provides computationally efficient testing procedures. Computing the MMD test statistic requires Monte Carlo (MC) integration. If $m$ denotes the number of MC samples, the overall computational cost becomes $O(n^{2}+2nm+m^{2})$ [gretton2012kernel]. When $m\ll n$, the cost is effectively quadratic; however, to control the MC error adequately, one often needs $m\gg n$, in which case the total complexity increases to $O(m^{2})$.
@@
A special class of IPM is KSD, which is also a special case of MMD [liu2016kernelized]. It can be shown to be characteristic under certain mild conditions [barp2024targeted]. In contrast to MMD, KSD does not require MC integration for any expectation. Leveraging the Stein’s identity, KSD test statistic can be computed with $O(n^{2})$ time complexity, although further acceleration is possible [jitkrittum2017linear] (see Appendix A.2 for details). Table 1 summarizes the computational complexity of the distance-based tests discussed above.
From a GoF testing perspective, compared to K-S, W-1 distances and MMD, KSD is significantly more efficient especially when it comes to multivariate settings. Therefore, KSD emerges as a natural choice for nonparametric score-based testing, especially in the multivariate setting. Nevertheless, for univariate cases, i.e. $d=1$, the higher computational cost of kernel-based methods may be justified by gains in test power, a topic we revisit in Appendix L.
3 Semiparametric KSD test
In the following sections, we will first review the preliminaries of KSD in Section 3.1 and then propose the main methodology in Section 3.2 with the investigation of its asymptotic behaviors in Section 3.3. We establish a calibration procedure based on parametric bootstrap and study its weak convergence in Section 3.4.
@@ 3.1 Preliminaries: kernelized Stein class and KSD
First introduced in the machine learning community, kernelized Stein discrepancy (KSD, [liu2016kernelized]) was proposed as a measure of model discrepancy in settings where the model likelihood is intractable. Its popularity is largely due to its computational efficiency (see Table 1), its desirable characteristic property that distinguishes between probability distributions, and, crucially, its ability to handle intractable models. We refer the reader to [anastasiou2023stein] for a comprehensive review of these developments. KSD belongs to the general class of Stein discrepancy (SD), which is defined via Stein operator $\mathcal{A}_{p}:\mathcal{F}_{d,d}\to\mathcal{F}_{d,1}$. Recalling $s_{p}=\nabla_{x}\log p(x)$ is the score function and Stein operator is defined as $\mathcal{A}_{p}g(x)=\sum_{j=1}^{d}\partial_{j}g_{j}(x)+g^{\top}(x)s_{p}(x)$, where $g\in\mathcal{F}_{d,d}$, and $g_{j}(x)$ denotes the $j$-th coordinate of $g(x)$. Under suitable regularity conditions, the following identity holds: $\mathbb{E}_{X\sim\mathcal{L}_{p}}[\mathcal{A}_{p}g(X)]=0$, as established, for example, in Proposition 1 of [gorham2015measuring]. This is the so-called Stein’s identity [stein1972bound, gorham2015measuring]. Using this identity, the SD between $\mathcal{L}_{p}$ and $\mathcal{L}_{q}$ can be defined as $\mathrm{SD}(\mathcal{L}_{p},\mathcal{L}_{q})=\sup_{g\in\overline{\mathcal{F}}_{d,d}}\left|\mathbb{E}_{p}[\mathcal{A}_{p}g(X)]-\mathbb{E}_{q}[\mathcal{A}_{p}g(X)]\right|$ for a compact set $\overline{\mathcal{F}}_{d,d}\subseteq\mathcal{F}_{d,d}$ with $\mathbb{E}_{p}[\mathcal{A}_{p}g(X)]=0$ by Stein’s identity. If we take $\mathcal{F}_{d,d}=\mathcal{H}^{d}$, the product space of a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, and $\overline{\mathcal{F}}_{d,d}$ to be the unit ball in $\mathcal{H}^{d}$, then the resulting SD is called the kernelized Stein discrepancy (KSD):
@@
We refer readers who are unfamiliar with RKHS to [kanagawa2018gaussian] for a detailed introduction and review. If we further define the function class to be the kernelized Stein class, $\mathcal{F}_{\mathrm{KSD}}=\{\mathcal{A}_{p}g:\|g\|_{\mathcal{H}^{d}}\leq 1\}$, then definition (6) fits the framework of the IPM in (4):
@@
with $\mathbb{E}_{p}[f(X)]=0$ for any $f\in\mathcal{F}_{\mathrm{KSD}}$. Furthermore, one can show under mild conditions on the kernel function, the KSD is characteristic and we refer interested readers to Appendix A.4 for more details.
@@ 3.2 SKSD test with general nuisance estimators
where the explicit form of $h$ is provided in Appendix A.1. The alternative formulation (7) enables a more computationally efficient estimate of KSD, as it only requires evaluating the expectation of a bivariate function $h$. A nice property of the SKSD test statistic is that it does not rely on the likelihood function of the model $\mathcal{L}_{\hat{\theta}_{n}}$. In fact, the function $h$ only involves the score function $s_{p_{\hat{\theta}_{n}}}$ under the model $\mathcal{L}_{\hat{\theta}_{n}}$ and kernel function $K$. Moreover, evaluating function $h$ does not require any Monte Carlo simulation and the V-statistic (7) can be computed in quadratic time (recall Table 1). Moreover, there are certain choices of kernel function that can reduce the complexity to even $O(n)$, such as linear kernel (See Appendix A.2 for more details).
Another appealing property of SKSD is the compatibility with general estimation procedure for $\hat{\theta}_{n}$. When $\hat{\theta}_{n}=\theta_{0}\in\Theta$, test statistic (SKSD) boils down to the KSD statistic [liu2016kernelized, chwialkowski2016kernel]. In most of statistical practice, however, the parameter $\theta_{0}$ is unknown and needs to be estimated from the data. The SKSD test statistic is designed to accommodate the general estimation procedure for $\theta_{0}$. We allow the estimation of $\hat{\theta}_{n}$ to be highly problem-specific and impose no restrictions on the estimation procedure, provided the resulting estimator satisfies a so-called uniform asymptotic linear estimate, formalized in the following definition.
@@
Throughout this paper, we will assume that $\hat{\theta}_{n}$ is a UALE with an estimation procedure $\mathcal{A}$, formalized in the following assumption.
@@
Definition 2 is a slightly stronger requirement than the classical asymptotic linear estimate (ALE) assumption [van2000asymptotic, Chapter 5,] because of the uniformity requirement, but still considerably general. Assumption 1 requires the estimator $\hat{\theta}_{n}$ to admit a limit $\theta_{0}$ in probability with asymptotic linear expansion uniformly over a neighborhood of the limit. In particular, classical parametric estimators such as M-estimator and minimum distance estimator [wolfowitz1957minimum] can be shown to be UALE under very mild regularity conditions. The formulation of these estimators can be found in Appendix A.3.
@@ 3.3 Consistency and asymptotic null distribution
It is worth noting that the limit of the test statistic $T_{n}(X,\hat{\theta}_{n})$ depends on the realization of the parameter $\theta_{0}$ in Assumption 1, which in turn can be related to the choice of estimation algorithm $\mathcal{E}$. In particular, under the null hypothesis any consistent estimator satisfies $\mathcal{L}=\mathcal{L}_{\theta_{0}}$, so the test statistic converges to zero as desired. However, under the alternative hypothesis, different estimation algorithms may lead to different limits $\theta_{0}$ even when the data are generated from the same distribution. For example, when $\mathcal{E}$ is empirical risk minimization (ERM) with respect to different loss functions, the corresponding $\mathcal{L}_{\theta_{0}}$ is the projection of the true distribution $\mathcal{L}$ onto the model class $\{\mathcal{L}_{\theta}:\theta\in\Theta\}$ under the discrepancy metrics induced by those losses. This freedom of choice can, however, substantially affect the power of the resulting test procedure. We investigate the impact of different estimation algorithms on power in Sections 5.1 and 5.3.
@@
$S^{*}(f,X,\theta)=\mathcal{A}_{\theta}f(X)+\big\langle\mathbb{E}_{X^{\prime}\sim P_{\theta}}\left[[\nabla_{\theta}s_{\theta}(X^{\prime})]^{\top}f(X^{\prime})\right],I(X,\theta)\big\rangle.$ (9)
To formally state the theorem, Let $K_{x}$ be feature map of $x$ in $\mathcal{H}$ (see Appendix A.1 for a rigorous definition) and $\mathbf{K}_{x}=K_{x}\bm{1}_{d}$, where $\bm{1}_{d}$ is the $d$-dimensional all-ones vector. The following theorem characterizes the asymptotic null distribution of the SKSD test statistic.
@@
Suppose $X_{1},\ldots,X_{n}$ are i.i.d. samples from $\mathcal{L}_{\theta_{0}}$ with $\theta_{0}\in\Theta$. Under Assumption 1 and regularity conditions in Appendix D, we have $nT_{n}$ converges in distribution to $W$ as $n\rightarrow\infty$, i.e.,
where $W\overset{d}{=}\sum_{u=1}^{\infty}\lambda_{u}Z_{u}^{2}$ with $Z_{u}$ to be independent standard normal variables and $\lambda_{u}$’s to be the eigenvalues of the operator $T_{\sigma}:\mathcal{H}^{d}\to\mathcal{F}_{d,1}$ defined as $T_{\sigma}(f)(x)=\mathbb{E}_{X\sim\mathbb{P}_{\theta_{0}}}[S^{*}(f,X,\theta_{0})S^{*}(\mathbf{K}_{x},X,\theta_{0})]$ for any $f\in\mathcal{H}^{d}$ and $x\in\mathcal{X}$.

### v2

@@ 2.2 Nonparametric score-based tests via characteristic IPMs
Tests based on IPMs can be powerful against nonparametric alternatives when sufficiently rich function classes $\mathcal{F}$ are used. The form and expressiveness of $\mathcal{F}$ fundamentally determine the capacity of the aggregated score statistic to capture discrepancies between the null and alternative. This insight has motivated a rich body of literature on hypothesis testing using IPMs [42, 86, 83, 63, 73]. In particular, we define an IPM is characteristic if
@@
Examples of popular characteristic IPMs and related divergences include $\{f:f=\mathbbm{1}(A),A\subseteq\mathcal{X}\}$, Kolmogorov-Smirnov (K-S) distance [85]; $\{f:|f(x)-f(y)|/\|x-y\|_{2}\leq 1\}$, Wasserstein-1 (W-1) distance [93]; a reproducing kernel Hilbert space (RKHS), maximum mean discrepancy [43]. Through the equivalence in (4), tests based on these IPMs can be interpreted as nonparametric score-based tests under the exponentially tilted models (2), provided that the IPM is characteristic. We now formally introduce the nonparametric score-based tests.
@@
According to the definition, the simplest way to verify that a score-based test is nonparametric is to evaluate if the corresponding IPM is characteristic, as per definition (5). It is well-known that the K-S and W-1 distances are characteristic under mild assumptions. Similarly, MMD is characteristic when appropriate kernels and regularity conditions are satisfied [42, 68]. Although the resulting test is characteristic when $\mathcal{F}$ is chosen to be sufficiently rich and complex, significant computational challenges arise for nonparametric score-based tests. To illustrate this, we now discuss the computational complexity of test statistics induced by several popular IPMs.
@@
In the univariate case ($d=1$), the classical K-S distance between an empirical distribution and a known distribution can be computed efficiently using sorting algorithms, yielding a computational complexity of $O(n\log n)$. However, in the bivariate case ($d=2$), the worst-case computational complexity increases to $O(n^{2})$ [74, 32]. When $d>2$, the K-S distance becomes NP-hard to compute [38].
@@
Similarly, when $d=1$, the W-1 distance can be computed efficiently in $O(n\log n)$ time via sorting. For $d>1$, however, the problem can be formulated as an assignment problem through discretization, and the computational complexity generally grows to $O(n^{3})$ [59], although approximation algorithms exist to improve efficiency [3].
@@
Computing the MMD test statistic requires numerical integration. Take Monte Carlo (MC) as a means for integration. If $m$ denotes the number of MC samples, the overall computational cost becomes $O(n^{2}+m^{2})$ [42]. When $m\ll n$, the cost is effectively quadratic; however, to control the MC error adequately, one needs $m\gg n$, in which case the total complexity increases to $O(m^{2})$.
@@
A special class of IPM is KSD, which is also a special case of MMD [63]. It can be shown to be characteristic under certain mild conditions [11]. In contrast to MMD, KSD does not require MC integration for any expectation. Leveraging the Stein’s identity (see Section 3.1 for details), KSD test statistic can be computed with $O(n^{2})$ time complexity, although further acceleration is possible [56] (see Appendix A.2 for details).
Table 1 summarizes the computational complexity of the distance-based tests discussed above. From a practical standpoint, KSD is substantially more efficient than K–S, W-1, and MMD, particularly in multivariate settings where the latter methods become computationally more demanding or even infeasible. As a result, KSD emerges as a natural and scalable choice for nonparametric score-based testing in moderate to high dimensions. In contrast, in the univariate case ($d=1$), the additional computational cost of kernel-based methods may be offset by gains in test power, an issue we revisit through simulation studies in Appendix L.
3 Semiparametric KSD test framework
In the following sections, we will first present the preliminaries of KSD in Section 3.1 and then propose the main methodology in Section 3.2 with the investigation of its asymptotic behaviors in Section 3.3. We establish a generic calibration procedure based on parametric bootstrap and study its weak convergence in Section 3.4.
@@ 3.1 Preliminaries: kernelized Stein class and KSD
First introduced in the machine learning community, kernelized Stein discrepancy (KSD, [63]) was proposed as a measure of model discrepancy in settings where the model likelihood is intractable. Its popularity is largely due to its computational efficiency (see Table 1), its desirable characteristic property that distinguishes between probability distributions (see Proposition 1), and, crucially, its ability to handle intractable models. We refer the reader to [4] for a comprehensive review of these developments.
KSD belongs to the general class of Stein discrepancy (SD), which is defined via Stein operator $\mathcal{A}_{p}:\mathcal{F}_{d,d}\to\mathcal{F}_{d,1}$. Recalling $s_{p}=\nabla_{x}\log p(x)$ is the score function and Stein operator is defined as $\mathcal{A}_{p}g(x)=\sum_{j=1}^{d}\partial_{j}g_{j}(x)+g^{\top}(x)s_{p}(x)$, where $g\in\mathcal{F}_{d,d}$, and $g_{j}(x)$ denotes the $j$-th coordinate of $g(x)$. Under suitable regularity conditions, the following identity holds: $\mathbb{E}_{X\sim\mathcal{L}_{p}}[\mathcal{A}_{p}g(X)]=0$, as established, for example, in Proposition 1 of [40]. This is the so-called Stein’s identity [87]. Using this identity, the SD between $\mathcal{L}_{p}$ and $\mathcal{L}_{q}$ can be defined as $\mathrm{SD}(\mathcal{L}_{p},\mathcal{L}_{q})=\sup_{g\in\overline{\mathcal{F}}_{d,d}}\left|\mathbb{E}_{p}[\mathcal{A}_{p}g(X)]-\mathbb{E}_{q}[\mathcal{A}_{p}g(X)]\right|$ for a compact set $\overline{\mathcal{F}}_{d,d}\subseteq\mathcal{F}_{d,d}$ with $\mathbb{E}_{p}[\mathcal{A}_{p}g(X)]=0$ by Stein’s identity. If we take $\mathcal{F}_{d,d}=\mathcal{H}^{d}$, the product space of a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, and $\overline{\mathcal{F}}_{d,d}$ to be the unit ball in $\mathcal{H}^{d}$, then the resulting SD is called the kernelized Stein discrepancy (KSD):
@@
We refer readers who are unfamiliar with RKHS to [57] for a detailed introduction and review. If we further define the function class to be the kernelized Stein class, $\mathcal{F}_{\mathrm{KSD}}=\{\mathcal{A}_{p}g:\|g\|_{\mathcal{H}^{d}}\leq 1\}$, then definition (6) fits the framework of the IPM in (4) by setting $\mathcal{F}=\mathcal{F}_{\textnormal{KSD}}$:
@@
with $\mathbb{E}_{p}[f(X)]=0$ for any $f\in\mathcal{F}_{\mathrm{KSD}}$. Furthermore, one can show under mild conditions on the kernel function, the KSD is characteristic (see Appendix A.4 for more details).
@@ 3.2 SKSD test with general nuisance estimators
where the explicit form of $h$ is provided in Appendix A.1. The alternative formulation (7) enables a computationally more efficient estimate of $\mathrm{KSD}^{2}$, as it only requires evaluating the expectation of a bivariate function $h$. A nice property of the SKSD test statistic is that it does not rely on the likelihood function of the model $\mathcal{L}_{\hat{\theta}_{n}}$. In fact, the function $h$ only involves the score function $s_{p_{\hat{\theta}_{n}}}$ under the model $\mathcal{L}_{\hat{\theta}_{n}}$ and kernel function $K$. Moreover, evaluating function $h$ does not require any Monte Carlo simulation and the V-statistic (7) can be computed in quadratic time (recalling Table 1). Moreover, there are certain choices of kernel function that can reduce the complexity to even $O(n)$, such as linear kernel (See Appendix A.2 for more details).
Another appealing property of SKSD is the compatibility with general estimation procedure for $\hat{\theta}_{n}$. We allow the estimation of $\hat{\theta}_{n}$ to be highly problem-specific and impose no restrictions on the estimation procedure, provided the resulting estimator satisfies a so-called uniform asymptotic linear estimate, formalized in the following definition.
@@
Definition 2 is a slightly stronger requirement than the classical asymptotic linear estimate (ALE) assumption [91, Chapter 5,] because of the uniformity requirement, but still considerably general. Throughout this paper, we will assume that $\hat{\theta}_{n}$ is a UALE with an estimation procedure $\mathcal{A}$, formalized in the following assumption.
@@
We make the following remark on the presented assumption.
Remark 1 (Broad applicability of Assumption 1).
Assumption 1 requires the estimator $\hat{\theta}_{n}$ to admit a limit $\theta_{0}$ in probability with asymptotic linear expansion uniformly over a neighborhood of the limit. In particular, classical parametric estimators such as M-estimator and minimum distance estimator [98] can be shown to be UALE under very mild regularity conditions. The formulation of these estimators can be found in Appendix A.3. When $\hat{\theta}_{n}=\theta_{0}\in\Theta$, test statistic (SKSD) boils down to the $\mathrm{KSD}^{2}$ statistic for point-null hypothesis [63, 23]. In most of statistical practice, however, the parameter $\theta_{0}$ is unknown and needs to be estimated from the data. The SKSD test statistic is designed to accommodate the general estimation procedure for $\theta_{0}$.
@@ 3.3 Consistency and asymptotic null distribution
It is worth noting that the limit of the test statistic $T_{n}(X,\hat{\theta}_{n})$ depends on the realization of the parameter $\theta_{0}$ in Assumption 1, which in turn can be related to the choice of estimation algorithm $\mathcal{E}$. In particular, under the null hypothesis any consistent estimator satisfies $\mathcal{L}=\mathcal{L}_{\theta_{0}}$, so the test statistic converges to zero as desired. However, under the alternative hypothesis, different estimation algorithms may lead to different limits $\theta_{0}$ even when the data are generated from the same distribution. For example, when $\mathcal{E}$ is empirical risk minimization (ERM, [92]) with respect to different loss functions, the corresponding $\mathcal{L}_{\theta_{0}}$ is the projection of the true distribution $\mathcal{L}$ onto the model class $\{\mathcal{L}_{\theta}:\theta\in\Theta\}$ under the discrepancy metrics induced by those losses. This freedom of choice can, however, substantially affect the power of the resulting test procedure. We investigate the impact of different estimation algorithms on power in Sections 5.1 and 5.3.
@@
$S^{*}(f,X,\theta)=\mathcal{A}_{\theta}f(X)+\big\langle\mathbb{E}_{X^{\prime}\sim\mathcal{L}_{\theta}}\left[[\nabla_{\theta}s_{\theta}(X^{\prime})]^{\top}f(X^{\prime})\right],I(X,\theta)\big\rangle.$ (9)
To formally state the theorem, Let $K_{x}$ be feature map of $x$ in $\mathcal{H}$ (see Appendix A.1 for a rigorous definition) and $\mathbf{K}_{x}\equiv K_{x}\bm{1}_{d}$, where $\bm{1}_{d}$ is the $d$-dimensional all-ones vector. The following theorem characterizes the asymptotic null distribution of the SKSD test statistic.
@@
Suppose $X_{1},\ldots,X_{n}$ are i.i.d. samples from $\mathcal{L}_{\theta_{0}}$ with $\theta_{0}\in\Theta$. Define $W\overset{d}{=}\sum_{u=1}^{\infty}\lambda_{u}Z_{u}^{2}$ with $Z_{u}$ to be independent standard normal variables and $\lambda_{u}$’s to be the eigenvalues of the operator $T_{\sigma}:\mathcal{H}^{d}\to\mathcal{F}_{d,1}$ defined as $T_{\sigma}(f)(x)=\mathbb{E}_{X\sim\mathcal{L}_{\theta_{0}}}[S^{*}(f,X,\theta_{0})S^{*}(\mathbf{K}_{x},X,\theta_{0})]$ for any $f\in\mathcal{H}^{d}$ and $x\in\mathcal{X}$. Then under Assumption 1 and regularity conditions in Appendix D, we have $nT_{n}$ converges in distribution to the random variable $W$, i.e.,
$\limsup_{n\rightarrow\infty}\left|\mathbb{P}_{H_{0}}[nT_{n}\leq t]-\mathbb{P}[W\leq t]\right|=0\quad\forall t\in\mathbb{R}.$